The potential at a point x due to some charges situated on the x axis is given by V=20/(x^2-4) volt.the electric field E at x =4 micrometre is given by
1
Expert's answer
2018-09-25T10:19:08-0400
V=20/(x^2-4).
E=-dV/dx i=-d/dx (20/(x^2-4))i=-20(-1/(x^2-4)^2 d/dx (x^2-4))i
E=20(1/(x^2-4)^2 2x)i=40x/(x^2-4)^2 i
The electric field E at x =4 micrometer is given by
E(4)=40(4)/(4^2-4)^2 =(10/9 i) V/μm